9. A force of 60 dynes acts upon a mass of 15g a) What acceleration is imparted to the body, b) What velocity will the body acquire in 8s? c) What distance will the body cover in this 8s?


1
Expert's answer
2021-09-24T09:27:19-0400

Given:

"F=60 \rm \: dyn=0.0006\: N"

"m=15\:\rm g=0.015\: kg"


(a) the acceleration

"a=\frac{F}{m}=\frac{0.0006\:\rm N}{0.015\:\rm kg}=0.04\:\rm m/s^2"

(b) the final speed

"v_f=v_i+at=\rm 0\: m/s+0.04\: m/s^2*8\: s=0.32\: m/s"

(d) the distance

"d=\frac{v+v_0}{2}t=\rm \frac{0\: m/s+0.32\: m/s}{2}*8\: s=1.28\: m"